which of the following equations have the solution k=5? show your work


a) k + 5=10 b) 5k=25 c) k - 5=10 d) k/6=5 (k over 6)


1
Expert's answer
2021-05-02T07:54:25-0400

"\text{Evaluating each of the options one by one we have:}\\
\text{(a) } k+5 = 10 \quad \text{Subtrating $5$ from both sides we have;}\\
k = 10-5 = 5\\
k = 5\\
\text{(b) } 5k = 25 \quad \text{Dividing both sides by $5$ we have: }\\
k = \frac{25}{5} = 5\\
k = 5\\
\text{(c) } k -5 = 10 \quad \text{Adding $5$ to both sides we have: }\\
k = 10 +5\\
k = 15\\
\text{(d) } \frac{k}{6} = 5 \quad \text{ Multiplying both sides by $6$ we have: }\\
k = 5 \times 6 = 30\\
k = 30\\

\text{From above we observed that only equations (a) } k+5=10 \text{ and (b) } 5k = 25 \text{ are the only equations which have the solution }k = 5"